NUCLEAR SPIN OF H_{3}^{+} IN DIFFUSE MOLECULAR CLOUDS

\underline{KYLE N. CRABTREE}, NICK INDRIOLO, HOLGER KRECKEL, BRIAN A. TOM,^{a} BENJAMIN J. McCALL, Department of Chemistry, University of Illinois, Urbana, IL 61801, USA.

In diffuse molecular clouds (environments with high molecular fraction, but low CO abundance), the relative populations of the J = 0 (para) and J = 1 (ortho) rotational levels of H_{2} are often used as a measure of the cloud kinetic temperature, T_{01}. Typically, T_{01} is on the order of 50-70 K, but in similar environments, the excitation temperature T (H_{3}^{+}) derived from the (J, K) = (1,0) (ortho) and (1,1) (para) rotational levels of H_{3}^{+} is 20-40 K. We have extended the number of sight lines in which both T_{01} and T (H_{3}^{+}) have been measured from two to five, and in four of the five cases, the two temperatures are discrepant in the same cloud. Using a steady state chemical model based on rate coefficients calculated with a microcanonical statistical approach, we find that the discrepancy between T_{01} and T (H_{3}^{+}) likely arises from incomplete thermalization caused by competition between the thermalization reaction H_{3}^{+} + H_{2} → H_{2} + H_{3}^{+} and dissociative recombination of H_{3}^{+} with electrons.

aPresent Address: Department of Chemistry, United States Air Force Academy, Air Force Academy, CO 80840, USA